Use an integrating factor to reduce the equations to exact equations and find the general solution.

6, (xy-1) dx + xidy = 0
Answer #1

(xy4) dx tx? dy = 0 Jaz Xリー dy ニ0 T- XP dy dy =) х2Compare with Standard form of first order linear ODE y + Pox) y = QX) SP) dx Integrating factor Clof) = Incx) = 2C A IntegraNow Solu tion is Given by y.(IF) = SQ0X) (FF) dx %3D ə yox = x •2주 ( (- Consten =Ino)+C, orintefsaT y = Incx)+ C General 801